3. Given w = x2 + y2, where x = eat and y = (4+t)2, use the chain rule to determine du
Answer #1


w = 1 + y = f(say), r = e, y = (4+t)

dw Of or of oy dt дz aт ду ді (Chain rule)

            Ә(x° + ) де“ да + *) Ә(4 +t)? +- дет ду at

             = 2. * 2e2+ + 2y + 2(4 + t) = 4.re2t + 4y(4+t)

               = 4e244 + 4(4+t)(4+t), 1 = e&y= (4+t)

  = 4e4 + 4(4+t)                 

dw = 4(e4 +(4+t)) dt